if amy can paint a house in 6 hours, and pat can paint the same houe in 4, and brian can paint the house in 7 how long would it take all of them o paint the house?
Answer by oscargut(686)

amy can paint a house in 6 hours then in 1 hour amy paints 1/6 of the house
pat can paint the same house in 4 then in 1 hour pat paints 1/4 of the house
brian can paint the house in 7 then in 1 hour amy paints 1/7 of the house
so working together in 1 hour they paint 1/6+1/4+1/7 of the house
so in one hour they paint (14+21+12)/84=47/84 of the house
so 1 hour 47/84
x hour 1
then x=84/47
Answer: in 84/47 hours